# Enhancing the Accuracy of Spatio-Temporal Models for Wind Speed Prediction by Incorporating Bias-Corrected Crowdsourced Data

### v1

@@ 4 Spatio-Temporal Modelling with Crowdsourced Wind Observations
Our goal is to investigate the impact of incorporating crowdsourced weather stations in the spatial modelling of wind speeds, and to explore modelling strategies that account for potentially noisy data. We focus on predicting wind speeds at unobserved sites using a computationally efficient framework that enables the generation of long-term, hourly wind speed time series at high spatial resolution. The approach also supports real-time mapping of wind speeds across Ireland, making it suitable for both retrospective analysis and operational use.
@@ 4.1 Spatio-Temporal Models for Wind Speeds
To capture the time-of-day effect, we model a smooth diurnal trend using a cyclic random walk of order 1, following Gómez-Rubio, (2020). Let $t\in\{1,\dots,24\}$ represent the hour of the day. We include a latent effect $d(t)$ modelled as:
@@
The final model including covariates is given as:
@@
Table 2: Classification of weather stations and their assigned noise parameters. The station categories are discussed in further detail in Section 2.2.
@@ 4.3 Prior Specification
For the IGP model, there are seven hyperparameters: the spatial effective range $\phi$, the standard deviation of the spatial process $\sigma_{z}$, and five parameters controlling the nugget (independent noise) for each measurement group, as specified in Equation (21). The AR1 model contains an additional parameter, $\rho$, which controls the temporal correlation.
@@
For the $\rho$ parameter in the AR1 model, analysis shows wind speeds show a temporal trend, therefore our prior will assume strong temporal correlation:
@@ 4.4 Evaluation methods
[…]
@@ 4.5 Space-time Simulation Studies
Each model is fit using both the IGP and AR1 covariance structures for the latent field. To assess predictive accuracy, we perform a leave-one-out cross-validation at each of the 23 meteorological station locations. At each test site, the remaining stations are used to predict wind speeds at all 100 time points.
@@
Noise Level ($\sigma$) Reliable Data Only ($\hat{\sigma}_{\epsilon}$) All Data (Constant Params) ($\hat{\sigma}_{\epsilon}$) All Data (Different Variance) ($\hat{\sigma}_{\text{Met}}$, $\hat{\sigma}_{\text{PWS-1}}$,$\hat{\sigma}_{\text{PWS-2}}$)
@@
Another result of note is that, although the data were simulated from a separable space-time model, and the AR1 model accurately captures the covariance parameters, there is little difference in predictive accuracy between the AR1 and IGP models. We discuss this finding further in Appendix G. We also noticed that the IGP model was substantially faster to fit than the AR1. Therefore, in applications where the goal is to interpolate a historical time series over a long period, an IGP model may be sufficient. Nevertheless, we still examine the accuracy of both approaches on our real wind dataset.
@@ 5 Results
We apply our models to the combined dataset of meteorological stations and bias-corrected PWS. To evaluate the impact of covariates and the inclusion of PWS, we compare several models of increasing complexity under both IGP and AR1 covariance structures. We adopt a leave-one-station-out cross-validation (LOSOCV) strategy, in which one meteorological station is sequentially excluded from model training and used solely for evaluation. Each excluded station is predicted at all observed time points using the remaining stations, focusing on spatial interpolation rather than temporal forecasting. For each test station, we predict hourly wind speeds over a six-month period from June 2024 to November 2024 (4,392 time points). This validation is repeated for all 23 meteorological stations.
@@
2 Meteorological Stations only, with covariates 1.41 0.18
@@
5 Met and PWS (bias correction), with covariates, single nugget variance 1.32 0.17
6 Met and PWS (bias correction), with covariates, group specific variance 1.32 0.17
AR1 Process
7 Met and PWS (bias correction), with covariates, group specific variance 2.01 0.31
@@
Results in Table 6 show in the first place that the AR1 model provides worse accuracy compared to the case when independent replicates in time are used, with a RMSE of 2.01 versus 1.32. This indicates that the AR1 structure may not be suitable for the Irish wind speed data. We examine in Appendix G the potential reasons for this poor performance, and will focus in the following on the results from the independent replicates approach.
@@
The comparison of accuracy results between the model 4 and model 5 shows that including covariates, especially the mean derived from reanalysis data and distance to the sea, significantly improved model accuracy with RMSE decreasing from 1.52 to 1.32. Although other commonly used covariates, such as altitude or orography, were not explicitly included, we assume these are accounted for within the reanalysis output, which utilizes physical models of topography to simulate wind speeds.
Finally, the most accurate models (5-6) are the ones incorporating both bias-corrected stations and reanalysis-derived covariates for the mean function. Including the bias-corrected weather stations reduced the RMSE from 1.41 to 1.32, a 6.5% reduction, and decreased the CRPS from 0.19 to 0.17, representing a 10.5% improvement compared to using only meteorological stations. It is however surprising that both models perform similarly despite the more complex and allegedly more realistic structure of measurement errors in model 6. To investigate the reason for this similar level of performance, Table 7 and Table 8 display the hyperparameters of the two models fit in INLA. The credible intervals for the hyperparameters are notably narrow, likely due to the large volume of data. When the model was fitted to shorter time periods as an experiment, the intervals became noticeably wider, suggesting that data volume influences parameter uncertainty. See Appendix H for the effect the number of timepoints has on the confidence intervals of the spatial parameters.
@@
$\phi$ (km) 458 445 471
$\sigma_{\text{z}}$ 0.325 0.319 0.331
$\sigma_{\epsilon}$ 0.261 0.259 0.263
@@
$\phi$ (km) 456 443 469
$\sigma_{\text{z}}$ 0.325 0.319 0.331
$\sigma_{\text{Met}}$ 0.266 0.264 0.269
$\sigma_{\text{A}}$ 0.271 0.258 0.263
$\sigma_{\text{B}}$ 0.241 0.237 0.246
$\sigma_{\text{C}}$ 0.268 0.264 0.272
$\sigma_{\text{U}}$ 0.253 0.249 0.257
@@
Results in Table 8 show that in opposition with our initial hypothesis suggesting that lower wind grades should have higher observation noise, all groups show similar noise parameters. In addition, comparing Tables 7 and 8 shows the level of noise in each of the station groups is similar to the overall noise when a single parameter is used in model 5, which explains the similar performance of the two models. This result may be due to preprocessing steps such as bias correction and filtering out low-correlation sites, which results in corrected PWS data being of similar accuracy to meteorological stations, as well as ensuring PWS stations have similar levels of variability to meteorological stations.
To test the model’s ability to detect unreliable groups of wind measurements, we fit a version where the ‘unknown rating’ group (U) remains uncorrected (7 of 49 stations) and reiterated the comparison between models 5 and model 6 in terms of model accuracy (Table 9) and parameter estimates (Tables 10 and 11).
@@
5 Met and PWS (bias correction), with covariates, single nugget variance 1.56 0.21
6 Met and PWS (bias correction), with covariates, group-specific variance 1.36 0.18
@@
Effective Range (km) 458 445 471
$\sigma_{\text{z}}$ 0.30 0.29 0.31
$\sigma_{\epsilon}$ 0.44 0.43 0.44
@@
The results in Table 9 show that model 5, characterized by a single likelihood for measurement errors, is strongly impacted by noisy data. The measurement error parameter ($\sigma_{\epsilon}$) increases from 0.26 to 0.44 and accuracy metrics degrade significantly (RMSE increases from 1.32 to 1.56 and CRPS from 0.17 to 0.21). In contrast, the multiple likelihood approach of model 6 is subject to a limited performance decrease (RMSE increases from 1.32 to 1.36 and CRPS goes from 0.17 to 0.18). In addition, the noise parameter $\sigma_{U}$ inflates significantly in this setting (Table 11), effectively mitigating the influence of data from poor-quality stations.
@@
6 Met and PWS (bias correction), with covariates, group specific variance 1.32
@@
Results in Table 12 show that our statistical model offers an improvement over ERA5 in the 6-month test period, with a RMSE of 1.32 versus 1.37. While the improvement is small, and there are stations where ERA5 is more accurate, this approach offers the advantage of being available in real-time. This improves the potential for short-term forecasts, an important tool for grid management. In contrast, ERA5 is only available after a five-day delay. The statistical approach also provides robust uncertainty quantification, which is not readily available with reanalysis products.
@@ 6 Discussion
Our results underscore the importance of bias correction when incorporating crowdsourced data into predictive models. Including the bias-corrected weather stations led to an 7% reduction in error, while including non-bias-corrected stations resulted in a significant increase in error. When incorporating crowdsourced stations, our leave-one-out predictive accuracy was similar to, and in some cases slightly better than, predictions by ERA5 (4% decrease in error overall). As the number of PWS increase in the future we expect the performance of models to further improve.
@@
Future work will explore accuracy using an independent dataset, such as wind farm data, which could offer a better test of the model’s generalisability. As wind farms typically operate on a sub-hourly level, this may require aggregating data at finer temporal resolutions. Since many PWS provide sub-hourly data, they could support the generation of high-resolution wind time series – a type of output not available from reanalysis datasets. Another area of future research is to combine ERA5 and weather station observations into a single model. While we considered them as competing models in this study, numerous studies have investigated jointly modelling station observations and gridded weather models (e.g. Forlani et al.,, 2020; Paciorek,, 2012).
To simplify computations, we employed a stationary covariance structure. We also compared a spatial-only approach with a separable AR1 model; however, the spatial-only approach resulted in lower errors. This suggests that an AR1 structure does not adequately capture the space-time interaction in wind data. Although this approach does not fully represent the underlying process, previous studies have shown that stationary or separable covariance structures can significantly reduce computational costs with only a minor loss in accuracy.
Nonetheless, considerable research has focused on efficient non-stationary and non-separable models, such as nearest-neighbour Gaussian processes or SPDE approaches that model advection and diffusion in space and time to capture transport processes (Datta et al.,, 2016; Clarotto et al.,, 2024). Future research may explore implementing these advanced models to better capture non-stationary and non-separable behaviour. These models could also be used to study the impact of crowdsourced data on short-term forecasting and nowcasting of wind speeds.

### v2

@@ 4 Spatio-Temporal Modelling with Crowdsourced Wind Observations
Our goal is to investigate the impact of incorporating data from crowdsourced weather stations in the spatial modelling of wind speeds, and to explore modelling strategies that account for potentially noisy data. We focus on predicting wind speeds at unobserved sites using a computationally efficient framework that enables the generation of long-term, hourly wind speed time series at high spatial resolution. The approach also supports real-time mapping of wind speeds across Ireland, making it suitable for both retrospective analysis and operational use.
@@ 4.1 Spatio-Temporal Models for Wind Speeds
To capture the time-of-day effect, we model a smooth diurnal trend using a cyclic random walk of order 1, following Gómez-Rubio, 2020a . Let $t\in\{1,\dots,24\}$ represent the hour of the day. We include a latent effect $d(t)$ modelled as:
@@
Wind speeds in Ireland exhibit clear seasonal patterns, as well as longer-term variability associated with climate change (Leahy and McKeogh,, 2013; Nolan et al.,, 2012). In principle, these effects could be incorporated through a cyclic annual random walk for seasonal variation and an additional long-term temporal component to capture climate-driven trends. However, reliable estimation of these components typically requires several years of data in order to separate long-term trends from short-term fluctuations, whereas our study spans only six months. For this reason, we restrict temporal variation to the diurnal cycle. Seasonal variations within the study period will also be partially accounted for by the spatial random effects, as measured observations will reflect the prevailing weather conditions. The final model including covariates is given as:
@@
Table 6: Classification of weather stations and their assigned noise parameters. The station categories are discussed in further detail in Section $2.$
@@ 4.2 Parameter Inference: SPDE approach
Another advantage of INLA is that the posterior distribution of any missing values is obtained automatically as part of the model fitting process (Gómez-Rubio, 2020b, ). Although the meteorological stations in our dataset contain observations at all time points, the PWS data include gaps. While handling missing data is not a primary focus of our analysis, the model nonetheless returns posterior predictions for missing PWS values, conditional on the observations from all stations.
@@ 4.3 Prior Specification
For the IGP model, there are seven hyperparameters: the spatial effective range $\phi$, the standard deviation of the spatial process $\sigma_{z}$, and five parameters controlling the nugget (independent noise) for each measurement group, as specified in Equation (21). The AR(1) model contains an additional parameter, $\rho$, which controls the temporal correlation.
@@
For the $\rho$ parameter in the AR(1) model, analysis shows wind speeds show a temporal trend, therefore our prior will assume strong temporal correlation:
@@ 4.4 Evaluation methods
$\text{CRPS}=\frac{1}{N}\sum_{i=1}^{N}\int_{-\infty}^{\infty}\left(F_{i}(y)-\mathbb{I}(y\geq y_{\text{obs},i})\right)^{2}dy,$ (29)
@@ 4.5 Space-time Simulation Studies
Each model is fit using both the IGP and AR(1) covariance structures for the latent field. To assess predictive accuracy, we perform a leave-one-out cross-validation at each of the 23 meteorological station locations. At each test site, the remaining stations are used to predict wind speeds at all 100 time points.
@@
Noise Level ($\sigma$) Reliable Data Only ($\hat{\sigma}_{\epsilon}$) All Data (Constant Params) ($\hat{\sigma}_{\epsilon}$) All Data (Different Variance) ($\hat{\sigma}_{\text{Met}}$, $\hat{\sigma}_{\text{PWS-1}}$$\hat{\sigma}_{\text{PWS-2}}$)
@@
Another result of note is that, although the data were simulated from a separable space-time model, and the AR(1) model accurately captures the covariance parameters, there is little difference in predictive accuracy between the AR(1) and IGP models. We discuss this finding further in Appendix 5.3. We also noticed that the IGP model was substantially faster to fit than the AR(1). Therefore, in applications where the goal is to interpolate a historical time series over a long period, an IGP model may be sufficient. Nevertheless, we still examine the accuracy of both approaches on our real wind dataset.
@@ 5 Results
5.1 Spatial Prediction Results
We apply our models to the combined dataset of meteorological stations and bias-corrected PWS. To evaluate the impact of covariates and the inclusion of PWS, we compare several models of increasing complexity under both IGP and AR(1) covariance structures. We adopt a leave-one-station-out cross-validation (LOSOCV) strategy, in which one meteorological station is sequentially excluded from model training and used solely for evaluation. Each excluded station is predicted at all observed time points using the remaining stations, focusing on spatial interpolation rather than temporal forecasting. For each test station, we predict hourly wind speeds over a six-month period from June 2024 to November 2024 (4,392 time points). This validation is repeated for all 23 meteorological stations.
@@ 5.1 Spatial Prediction Results
2 Meteorological Stations only, with covariates 1.35 0.18
@@
5 Met and PWS (bias correction), with covariates, single nugget variance 1.28 0.16
6 Met and PWS (bias correction), with covariates, group specific variance 1.28 0.16
AR(1) Process
7 Met and PWS (bias correction), with covariates, group specific variance 1.90 0.30
@@
Results in Table 10 show in the first place that the AR(1) model provides worse accuracy compared to the case when independent replicates in time are used, with a RMSE of 1.9 versus 1.28. This indicates that the AR(1) structure may not be suitable for the Irish wind speed data. We examine in Appendix 5.3 the potential reasons for this poor performance, and will focus in the following on the results from the independent replicates approach.
@@
The comparison of accuracy results between the model 4 and model 5 shows that including covariates, especially the mean derived from reanalysis data and distance to the sea, significantly improved model accuracy with RMSE decreasing from 1.63 to 1.28. Although other commonly used covariates, such as altitude or orography, were not explicitly included, we assume these are accounted for within the reanalysis output, which utilizes physical models of topography to simulate wind speeds.
Finally, the most accurate models (5-6) are the ones incorporating both bias-corrected stations and reanalysis-derived covariates for the mean function. Including the bias-corrected weather stations reduced the RMSE from 1.35 to 1.28, a 5.2% reduction, and decreased the CRPS from 0.18 to 0.18, representing a 10.5% improvement compared to using only meteorological stations. It is however surprising that both models perform similarly despite the more complex and allegedly more realistic structure of measurement errors in model 6. To investigate the reason for this similar level of performance, Table 11 and Table 12 display the hyperparameters of the two models fit in INLA. The credible intervals for the hyperparameters are notably narrow, likely due to the large volume of data. When the model was fitted to shorter time periods as an experiment, the intervals became noticeably wider, suggesting that data volume influences parameter uncertainty. See Appendix G for the effect the number of timepoints has on the confidence intervals of the spatial parameters.
@@
$\phi$ (km) 526 513 535
$\sigma_{\text{z}}$ 0.326 0.321 0.330
$\sigma_{\epsilon}$ 0.282 0.281 0.284
@@
$\phi$ (km) 491 482 500
$\sigma_{\text{z}}$ 0.336 0.331 0.340
$\sigma_{\text{Met}}$ 0.259 0.258 0.261
$\sigma_{\text{A}}$ 0.285 0.277 0.294
$\sigma_{\text{B}}$ 0.297 0.294 0.300
$\sigma_{\text{C}}$ 0.315 0.312 0.317
$\sigma_{\text{U}}$ 0.274 0.271 0.277
@@
Results in Table 12 show that in opposition with our initial hypothesis suggesting that lower wind grades should have higher observation noise, all groups show similar noise parameters. In addition, comparing Tables 11 and 12 shows the level of noise in each of the station groups is similar to the overall noise when a single parameter is used in model 5, which explains the similar performance of the two models. This result may be due to preprocessing steps such as bias correction and filtering out low-correlation sites, which results in corrected PWS data being of similar accuracy to meteorological stations, as well as ensuring PWS stations have similar levels of variability to meteorological stations.
To test the model’s ability to detect unreliable groups of wind measurements, we fit a version where the ‘unknown rating’ group (U) remains uncorrected (8 of 49 stations) and reiterated the comparison between models 5 and model 6 in terms of model accuracy (Table 13) and parameter estimates (Tables 14 and 15).
@@
5 Met and PWS (bias correction), with covariates, single nugget variance 1.53 0.20
6 Met and PWS (bias correction), with covariates, group-specific variance 1.32 0.18
@@
Effective Range (km) 581 570 596
$\sigma_{\text{z}}$ 0.31 0.30 0.31
$\sigma_{\epsilon}$ 0.44 0.44 0.45
@@
The results in Table 13 show that model 5, characterized by a single likelihood for measurement errors, is strongly impacted by noisy data. The measurement error parameter ($\sigma_{\epsilon}$) increases from 0.26 to 0.44 and accuracy metrics degrade significantly (RMSE increases from 1.28 to 1.53 and CRPS from 0.16 to 0.20). In contrast, the multiple likelihood approach of model 6 is subject to a limited performance decrease (RMSE increases from 1.28 to 1.32 and CRPS goes from 0.16 to 0.17). In addition, the noise parameter $\sigma_{U}$ inflates significantly in this setting (Table 15), effectively mitigating the influence of data from poor-quality stations.
@@
6 Met and PWS (bias correction), with covariates, group specific variance 1.28
@@
Results in Table 16 show that our statistical model offers an improvement over ERA5 in the 6-month test period, with a RMSE of 1.28 versus 1.37. While the improvement is small (6.5%), and there are stations where ERA5 is more accurate, this approach offers the advantage of being available in real-time. This improves the potential for short-term forecasts, an important tool for grid management. In contrast, ERA5 is only available after a five-day delay. The statistical approach also provides robust uncertainty quantification, which is not readily available with reanalysis products.
@@
5.2 Performance under Extreme Wind Conditions
We also assess how accurately our model predicts extreme wind speeds. In the context of renewable energy, extreme wind speeds lead to wind curtailment and can also lead to damage to wind turbines(Hering and Genton,, 2010). Outside of renewable energy applications, extreme wind speeds can potentially damage built structures and pose a risk to life. To assess the accuracy of predicting extreme wind speeds we compare the RMSE, bias (predicted - actual) and correlation of wind speeds for the top 1%, 2.5% and 5% of wind speeds. Table 17 contains the metrics for our final model 6, and compared to ERA5.
Percentile Model RMSE Bias Pearson Correlation
5% Met + PWS (bias corrected), covariates, group–specific variance 2.43 -1.78 0.82
ERA5 2.35 -1.27 0.73
2.5% Met + PWS (bias corrected), covariates, group–specific variance 2.79 -2.15 0.82
ERA5 2.60 -1.50 0.74
1% Met + PWS (bias corrected), covariates, group–specific variance 3.20 -2.54 0.83
ERA5 2.89 -1.80 0.75
Table 17: Predictive performance under extreme wind conditions. Metrics reported for the top 5%, 2.5% and 1% of observed wind speeds.
Across all thresholds, the model tends to underpredict the highest wind speeds, as seen by the negative bias values. Analysis of the underlying data shows that extreme wind speeds in Ireland are often highly localised, especially at exposed coastal stations, and are not typically observed at nearby inland stations. Because our model relies on spatial information borrowed from neighbouring sites, these isolated extremes are more difficult to capture. In contrast ERA5, although it exhibits lower overall accuracy, performs slightly better in the upper tail. ERA5 gains from its physics based model, which capture storm conditions, as well as modelling wind fields over the surrounding ocean. As more observations are included (e.g., top 5% rather than top 1%), the performance gap narrows, and our model exhibits consistently higher correlation with the observed wind speeds.
5.3 Performance of AR(1) versus Independent Gaussian Processes
Across both the simulation study and the wind speed dataset is that the independent Gaussian process (IGP) model performs comparably to, or even outperforms the AR(1) model, despite the latter’s theoretical advantage in capturing temporal correlation. In the simulation study, this result can be explained by the separable spatio-temporal process used for data generation. For a separable process, the value at an unobserved location conditionally depends only on other locations at the same time point, and its own past or future values at the same location. Since the test case involves predictions at a completely unobserved location, the temporal component of the AR(1) model provides no additional information, reducing its predictive advantage over the purely spatial IGP model, see Table 8.
For the wind speed data, the estimated AR(1) correlation parameter is very close to 1 ($\rho\approx 0.98$). Notably, the spatial range parameter in the AR(1) model is smaller than that of the IGP model, indicating that the AR(1) model attributes most of the observed dependence to temporal persistence rather than spatial structure (See Table 18). Consequently, predictions at unobserved locations often tend toward the station-specific mean, as the reduced spatial range limits borrowing of information across stations. This behaviour likely explains the IGP’s superior performance—its larger spatial range allows it to leverage cross-station correlations more effectively.
The limited range may arise from how the AR(1) parameters are inferred. The spatio-temporal process at time $t$ is expressed as
$z_{s,t}=\rho z_{s,t-1}+\xi_{s,t},$
$\xi_{s,t}$ represents a spatially correlated innovation term between $t$ and $t+1$, which is used to infer spatial hyperparameters. Empirically, the hourly innovations in the wind data show much lower spatial correlation than expected under a true AR(1) process. In Figure 17 the correlation of hourly changes against distance is compared in both the wind dataset and the simulated data from a true AR(1) process, demonstrating how hourly changes in the wind dataset are less correlated than expected in simulated data. This mismatch likely leads the AR(1) model to underestimate the spatial range and attribute most of the structure to temporal dependence, thereby reducing its predictive performance relative to the IGP.
Parameter Parameter Value - IGP Parameter Value - AR(1)
Effective Range (km) 491 103
$\sigma_{\text{z}}$ 0.336 0.57
$\sigma_{\text{Met}}$ 0.259 0.123
$\sigma_{\text{A}}$ 0.285 0.221
$\sigma_{\text{B}}$ 0.297 0.158
$\sigma_{\text{C}}$ 0.315 0.240
$\sigma_{\text{U}}$ 0.274 0.162
$\rho$ 0.97
Table 18: Estimated parameters for Model 6 under both the IGP and AR(1) covariance structure.
(a) (a)
(b) (b)
Figure 17: A plot of empirical correlation against distance (in kilometres) for the hourly increments for (a) the wind speed dataset and (b) the simulated data from a true AR(1) process in Section 4.5.
@@ 6 Discussion
Our results underscore the importance of bias correction when incorporating crowdsourced data into predictive models. Including the bias-corrected weather stations led to an 5% reduction in error, while including non-bias-corrected stations resulted in a significant increase in error. When incorporating crowdsourced stations, our leave-one-out predictive accuracy was similar to, and in some cases slightly better than, predictions by ERA5 (6.5% decrease in error overall). As the number of PWS increase in the future we expect the performance of models to further improve.
@@
Future work will explore accuracy using an independent dataset, such as wind farm data, which could offer a better test of the model’s generalisability. As wind farms typically operate on a sub-hourly level, this may require aggregating data at finer temporal resolutions. Since many PWS provide sub-hourly data, they could support the generation of high-resolution wind time series, a type of output not available from reanalysis datasets. Another area of future research is to combine reanalysis data such as ERA5 and weather station observations into a single model. While we considered them as competing models in this study, numerous studies have investigated jointly modelling station observations and gridded weather models (e.g. Forlani et al.,, 2020; Paciorek,, 2012). In forecasting applications, integrating observations with NWP has been shown to yield superior predictive accuracy compared with using NWP alone (Chen et al., 2021b, ).
To simplify computations, we employed a stationary covariance structure. We also compared a spatial-only approach with a separable AR(1) model; however, the spatial-only approach resulted in lower errors. This suggests that an AR(1) structure does not adequately capture the space-time interaction in wind data. Although this approach does not fully represent the underlying process, previous studies have shown that stationary or separable covariance structures can significantly reduce computational costs with only a minor loss in accuracy.
Nonetheless, considerable research has focused on efficient non-stationary and non-separable models, such as nearest-neighbour Gaussian processes or SPDE approaches that model advection and diffusion in space and time to capture transport processes (Datta et al.,, 2016; Clarotto et al.,, 2024). Future research may explore implementing these advanced models to better capture non-stationary and non-separable behaviour, which could improve short-term forecasts.